Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Here,  n=10,  p=12,  q=1−12=12           (for  true/false  questions)and  r≥8  i.e.,  8, 9, 10∴  P(X≥8)=P(x=8)+P(x=9)+P(x=10)                          =C810(12)8(12)2+C910(12)9(12)+C1010(12)10(12)0                          =45.(12)10+10.(12)10+(12)10=(12)10(45+10+1)                          =56×11024=7128Hence,  the  correct  option  is  (b).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Probability  of  getting  Queen=452So,  the  required  probability=P(Queen).P(Queen)????????????????????????                                               =452.452=113.113             (with  replacement)Hence,  the  correct  option  is  (a).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

We  know  that  for  a  Binomial  distribution,   the  outcomes  must  not  be  dependent  on  each  other.Hence,   the  correct  option  is   (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Let  E1  be  the  event  showing  the  sum  of  the  numbers  on  the  two  dice  was  less  than  6∴            E1={(1,1),(1,2),(2,1),(1,3),(3,1),(1,4),(4,1),(2,2),(2,3),(3,2)}        n(E1)=10.  E2  be  the  event  that  the  sum  of  the  numbers  is  3.∴            E2={(1,2),(2,1)}        n(E2)=2  and  n(E1∩E2)=2∴P(E2/E1)=n(E1∩E2)n(E1)=210=15Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

  probability=P (dead).P (dead)? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ?                                                =38.27=328Hence,   the  correct  option  is   (d).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Probability  of  drawing  2  green  and  1  blue  balls                 =P(G).P(G).P(B)+P(G).P(B).P(G)+P(B).P(G).P(G)                 =38.27.26+38.27.26+28.37.26=12336+12336+12336=36336=328Hence,  the  correct  option  is  (a).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Let  E1  be  the  event  of  getting  even  number  on  the  die.  E2  be  the  event  of  selecting  a  spade  card.∴    P(E1)=36=12  and  P(E2)=1352=14So,           P(E1∩E2)=P(E1).P(E2)=12.14=18Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Let  G  denotes  the  girl  and  B  denotes  the  boy  of  the  given  family.So,  n(S)={(BGG),(GBG),(GGB),(GBB),(BGB),(BBG),(BBB),(GGG)}Let  E1  be  the  event  that  the  family  has  atleast  one  girl∴            E1={(BGG),(GBG),(GGB),(GBB),(BGB),(BBG),(GGG)}.  E2  be  the  event  that  the  eldest  child  is  a  girl.∴            E2={(GBG),(GGB),(GBB),(GGG)}(E1∩E2)={(GBB),(GGB),(GBG),(GGG)}∴P(E2/E1)=P(E1∩E2)P(E1)=4/87/8=47Hence,  the  correct  option  is  (d).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Given  that:  P(A)=0.4,  P(B)=0.3,  P(C)=0.2Also,  P(A¯)=1−0.4=0.6,  P(B¯)=1−0.3=0.7,  P(C¯)=1−0.2=0.8∴  Probabilities of  two  hits                 =P(A).P(B).P(C¯)+P(A).P(B¯).P(C)+P(A¯).P(B).P(C)                 =0.4×0.3×0.8+0.4×0.7×0.2+0.6×0.3×0.2                 =0.096+0.056+0.036=0.188Hence,  the  correct  option  is  (b).

Thumbs Up IconUpvote Thumbs Down Icon